A survey is being planned to determine the average amount of time preschool children watch television. A pilot survey indicated that the mean time per week is 12 hours, with a standard deviation of 3 hours. It is desired to estimate the mean viewing time within one-quarter hour. The 0.95 degree of confidence is to be used. How do you calculate how many preschool children should be surveyed? Why is it important to make sure your sample size is sufficient?
Answer #1

Population standard deviation, σ = 3

Margin of error, E = 1/4 = 0.25

Confidence interval, CL = 0.95

Significance level, α = 1-CL = 0.05

Critical value, z = NORM.S.INV(0.05/2) = 1.96

Sample size, n = (z * σ / E)² = (1.96 * 3 / 0.25)² = 553.17 = 553